# Bidirectional Decoding: Improving Action Chunking via Closed-Loop Resampling

### v3

@@ Abstract
Predicting and executing a sequence of actions without intermediate replanning, known as action chunking, is increasingly used in robot learning from human demonstrations. Yet, its reported effects on the learned policy are inconsistent: some studies find it crucial for achieving strong results, while others observe decreased performance. In this paper, we first dissect how action chunking impacts the divergence between a learner and a demonstrator. We find that action chunking allows the learner to better capture the temporal dependencies in demonstrations but at the cost of reduced reactivity in stochastic environments. To address this tradeoff, we propose Bidirectional Decoding (BID), a test-time inference algorithm that bridges action chunking with closed-loop operations. BID samples multiple predictions at each time step and searches for the optimal one based on two criteria: (i) backward coherence, which favors samples that align with previous decisions; (ii) forward contrast, which seeks samples of high likelihood for future plans. By coupling decisions within and across action chunks, BID promotes consistency over time while maintaining reactivity to unexpected changes. Experimental results show that BID boosts the performance of two state-of-the-art generative policies across seven simulation benchmarks and two real-world tasks. Code and videos are available at https://bid-robot.github.io.
@@ 1 Introduction
In this paper, we first dissect the influence of action chunking by examining the divergence between learned policies and human demonstrations. We find that, when a policy is built with limited context length – little or no history is used as input for robustness or efficiency (Mandlekar et al., 2020; Bharadhwaj et al., 2023b; Brohan et al., 2023a; b; Shi et al., 2023; Collaboration, 2023) – increasing the length of action chunks allows for implicit conditioning on more past actions, thereby improving its ability to capture the temporal dependencies inherent in demonstrations. However, this advantage comes at the cost of reduced access to recent state observations, which can be crucial for reacting to unexpected changes in stochastic environments, such as action noise or object motions. This tradeoff raises a crucial question: How can we preserve the strengths of action chunking in long-term consistency without suffering from its limitations in short-term reactivity?
To this end, we introduce Bidirectional Decoding (BID), an inference algorithm that bridges action chunking with closed-loop operations. Our main idea is to sample multiple predictions at each time step and search for the most desirable one. Specifically, BID operates on two decoding criteria: (i) backward coherence, which favors samples that are close to the sequence selected at the previous step; (ii) forward contrast, which favors samples that are close to the output of a stronger policy and distant from those of a weaker one. As illustrated in Fig. 1, BID updates the chunk of future actions based on the previous strategy, promoting temporal consistency over extended periods while remaining reactive to unexpected changes.
@@ 3.1 Preliminaries
Notably, Zhao et al. (2023b) hypothesize that action chunking mitigates compounding errors, but it is unclear why this would hold when deviations within a chunk cannot be corrected before replanning. Alternatively, Chi et al. (2023) draws parallels to model predictive control (MPC) (Borrelli et al., 2017; Löfberg, 2012), emphasizing its role in improving consistency and planning. However, unlike MPC, which typically uses short action horizons (e.g., $h=1$), recent imitation learning methods employ much longer action horizons, such as a substantial portion of the predicted sequence (e.g., $h=8$ in Chi et al. (2023)) or even the entire prediction length (e.g., $h\geq 50$ in Zhao et al. (2023b); Black et al. (2024)). The lack of a clear understanding of action chunking hinders its effective use across tasks and environments. In the next sections, we seek to explicitly analyze the strengths and weaknesses of action chunking, with particular attention to the choice of action horizon $h$, in order to offer practical guidance and to inspire the design of improved algorithms.
@@ 3.2 Analysis
Additionally, since a $(c,h)$-policy observes only a subset of the states that the expert is conditioned on, we assume that an optimal policy must reconstruct all missing information correctly:
Assumption 2. An optimal $\pi_{c,h}$ must infer the unobserved states based on the observed states and actions by modeling the transition dynamics $P(s_{t^{\prime}}\mid s_{t^{\prime}-1},a_{t^{\prime}-1})$ accurately for all time step $t^{\prime}$.
Under these assumptions, the divergence between a learned policy and an expert policy is attributed to two factors: (i) the importance of unobserved states in predicting the current action, and (ii) the difficulty of inferring unobserved states based on the available information.
To more clearly see the influence of action horizon on these factors, we next compare the performance of two policies that have the same context lengths but different action horizons, $\pi_{h}\coloneqq\pi_{(c,h)}(a_{t}|s_{t-h-c:t-h},a_{t-h:t-1})$ and $\pi_{h+d}\coloneqq\pi_{(c,h+d)}(a_{t}|s_{t-h-d-c:t-h-d},a_{t-h-d:t-1})$, where $d>0$ is the extended action horizon. As illustrated in Fig. 3, each policy has access to unique information that is unavailable to the other. $\pi_{h}$ observes some recent states, where $\pi_{h+d}$ is only aware of the executed actions. On the other hand, $\pi_{h+d}$ has access to some earlier states and actions, which precede all information available to $\pi_{h}$. We characterize the importance of observations as follows (formal definitions in Sec. D.1):
@@
Hence, we denote the observation advantage that $\pi_{h}$ gains from the observed recent states by $\alpha_{f}$ and the inference disadvantage it incurs from the earlier unobserved states by $\epsilon_{b}$, whereas $\pi_{h+d}$ conversely gains $\alpha_{b}$ but incurs $\epsilon_{f}$.
The difficulty of inferring each unobserved state hinges on both the relevant observations as well as the environmental stochasticity. We quantify this difficulty as follows:
Definition (Forward Inference). Let $P_{f}\coloneqq P(S_{t}=g_{t}|S_{t-1}=g_{t-1},a_{t-1})$ where $g_{t}$ and $g_{t-1}$ are the ground truth states in the deterministic environment at time $t$ and $t-1$, respectively. In deterministic environments, $P_{f}=1$, whereas in stochastic settings, $P_{f}$ is smaller.
Definition (Backward Inference). Let $P_{b}\coloneqq P(S_{t}=g_{t}|S_{t+1}=g_{t+1})$ where $g_{t}$ and $g_{t+1}$ are the ground truth states in the deterministic environment at time $t$ and $t+1$, respectively. Since $P_{b}$ is not conditioned on any action, it has higher entropy in general. In stochastic environments, $P_{b}$ is small.
Given that the forward inference is generally easier than the backward inference, the performances of $\pi_{h}$ and $\pi_{h+d}$ differ as follows (proofs are deferred to Appendix D):
@@
Let $\mathcal{L}$ be a non-linear convex function measuring the prediction error with respect to demonstrations. Let $\mathcal{S}^{+}\subset\{s_{t-k:t}\}$ be the states both the $(c,h)$ and the $(c,h+d)$ policies observe and let $\mathcal{S}^{-}\coloneqq\{s_{t-k:t}\}\setminus\mathcal{S}^{+}$. Similarly, let $\mathcal{P}^{+}\coloneqq\{a_{t-h:t-1}\}$ and $\mathcal{P}^{-}\coloneqq\{a_{t-h-d:t-h-1}\}$. Let $C\coloneqq\mathcal{P}^{+}\cup\mathcal{S}^{+}$, $G\coloneqq\{a_{t},z_{t-k:t}\}\cup\mathcal{P}^{-}\cup\mathcal{S}^{-}$. Then, we can bound the expected loss of the $(c,h+d)$-policy and the $(c,h)$-policy as:
$\alpha_{f}-\epsilon_{b}(1-P_{b}^{2d})\leq\min_{\pi_{h+d}}\mathbb{E}_{G}\left[\mathcal{L}(\pi_{h+d},\pi^{*})|C\right]-\min_{\pi_{h}}\mathbb{E}_{G}\left[\mathcal{L}(\pi_{h},\pi^{*})|C\right]\leq-\alpha_{b}+\epsilon_{f}(1-P_{f}^{2d}).$ (2)
@@
Remark 1.1. Eq. 2 provides a general comparison of the performance of the two policies. Intuitively, the advantage of each policy stems from the additional information it has access to (i.e. $\alpha_{f}$ for $\pi_{h}$ and $\alpha_{b}$ for $\pi_{h+d}$) while the disadvantage is bounded by the maximum divergence arising from inferring missing information incorrectly (i.e. $\epsilon_{b}(1-P_{b}^{2d})$ for $\pi_{h}$ and $\epsilon_{f}(1-P_{f}^{2d})$ for $\pi_{h+d}$).
@@
We next examine two specific environmental settings: deterministic and stochastic. In highly deterministic environments, while both policies need to infer the same number of unobserved states, $\pi_{h+d}$ benefits from conditioning on additional actions, which may significantly aid in inferring the corresponding states through its action chunk. If the maximum errors $\epsilon_{f}$ arising from inferring these states are bounded, $\pi_{h+d}$ becomes strictly advantageous:
Corollary 2 (Consistency in Deterministic Environments).
In a highly deterministic environment, if $a_{t}$ is influenced by at least one state in $\{s_{t-h-c-d:t-h-c-1}\}$ and $\epsilon_{f}$ is finite, then
$\min_{\pi_{h+d}}\mathbb{E}_{G}\left[\mathcal{L}(\pi_{h+d},\pi^{*})|C\right]<$ $\min_{\pi_{h}}\mathbb{E}_{G}\left[\mathcal{L}(\pi_{h},\pi^{*})|C\right].$ (3)
Conversely, in highly stochastic environments, inferring the unobserved states is challenging, regardless of whether the corresponding actions are known. On the other hand, the recent states are likely more important than earlier states for predicting the current action $a_{t}$. As a result, $\pi_{h+d}$ becomes strictly more disadvantageous:
Corollary 3 (Reactivity in Stochastic Environments).
In a highly stochastic environment, if temporal dependency decreases over time, i.e., $\alpha_{f}>\epsilon_{b}$, then
$\min_{\pi_{h+d}}\mathbb{E}_{G}\left[\mathcal{L}(\pi_{h+d},\pi^{*})|C\right]>$ $\min_{\pi_{h}}\mathbb{E}_{G}\left[\mathcal{L}(\pi_{h},\pi^{*})|C\right].$ (4)
@@
In summary, depending on the experimental conditions, action chunking may have varying effects on the learned policy. On the one hand, it benefits the modeling of temporal dependencies in the demonstrations, due to extended access to actions executed at past time steps. On the other hand, it hinders reactions to unexpected states in stochastic environments, due to reduced access to state observations at recent time steps. As a result, there is no universally optimal choice of action horizon across all conditions. When both temporal dependencies and transition stochasticities are significant, tuning the action horizon entails an inherent trade-off between the two opposing factors. Moreover, stochasticitic changes (e.g., wind gusts or object drops) are often unpredictable in advance, raising a critical need to ensure the reactivity of a policy built with action chunking.
@@ 4 Method: Bidirectional Decoding
As analyzed above, action chunking improves long-term consistency but sacrifices short-term reactivity. In this section, we propose to address this tradeoff by bridging action chunking with closed-loop operations. We will first outline the general framework in Sec. 4.1 and then describe two specific criteria in Sec. 4.2.
@@ 4.2 Bidirectional Criteria
Figure 4: Illustration of bidirectional decoding.
Algorithm 1 Bidirectional Decoding
2:  Compute the backward loss $\mathcal{L}_{B}$ for each sample
4:  Compute the forward loss $\mathcal{L}_{F}$ for each sample
5:  Select $a^{*}\in\mathcal{A}$ that minimizes the total loss
6:  Update decision memory $\hat{a}\leftarrow a^{*}$
Fig. 4 illustrates the combined effects of the backward coherence and forward contrast criteria on sample selection. Since samples in $\mathcal{A}^{+}$ and $\mathcal{A}^{-}$ are not all subject to the same strategy, we trim each set by removing samples that deviate significantly from the previous decision. This is achieved by summing over the $K$ smallest distance values for in the positive and negative sets in Eq. 7. The full process of our decoding method is outlined in Algorithm 1. Since all steps in BID can be computed in parallel, the overall computational overhead remains modest on modern GPU devices.
4.3 Discussions
5 Experiments
@@ 5.2.1 Comparison with Existing Inference Methods
We evaluate each method for $100$ episodes and average the results across three random seeds. Please refer to Appendix B for implementation details.
@@ 5.3 Real-world Experiments with Dynamic Objects
Setup. We consider two pick-and-place tasks, where the target object undergoes unexpected movement during evaluation. In the dynamic placing task, a Franka Panda robot is required to deliver an object into a moving cup held by a human subject. In the dynamic picking task, a UR5 robot is required to grasp a moving cup pulled by a string and place it onto a nearby static saucer. In both tasks, we evaluate the performance of BID applied to pre-trained diffusion policies. For further experimental details, please refer to Sec. B.2.

### v4

# Bidirectional Decoding: Improving Action Chunking via Guided Test-Time Sampling

@@ Abstract
Predicting and executing a sequence of actions without intermediate replanning, known as action chunking, is increasingly used in robot learning from human demonstrations. Yet, its effects on the learned policy remain inconsistent: some studies find it crucial for achieving strong results, while others observe decreased performance. In this paper, we first dissect how action chunking impacts the divergence between a learner and a demonstrator. We find that action chunking allows the learner to better capture the temporal dependencies in demonstrations but at the cost of reduced reactivity to unexpected states. To address this tradeoff, we propose Bidirectional Decoding (BID), a test-time inference algorithm that bridges action chunking with closed-loop adaptation. At each timestep, BID samples multiple candidate predictions and searches for the optimal one based on two criteria: (i) backward coherence, which favors samples that align with previous decisions; (ii) forward contrast, which seeks samples of high likelihood for future plans. By coupling decisions within and across action chunks, BID promotes both long-term consistency and short-term reactivity. Experimental results show that our method boosts the performance of two state-of-the-art generative policies across seven simulation benchmarks and two real-world tasks. Videos and code are available at https://bid-robot.github.io.
@@ 1 Introduction
In this paper, we first dissect the influence of action chunking by examining the divergence between learned policies and human demonstrations. We find that, when a policy is built with limited context length – little or no history is used as input for robustness or efficiency (Mandlekar et al., 2020; Bharadhwaj et al., 2023b; Brohan et al., 2023a; b; Shi et al., 2023; Collaboration, 2023) – increasing the length of action chunks allows for implicit conditioning on more past actions, thereby improving its ability to capture the temporal dependencies inherent in demonstrations. However, this advantage comes at the cost of reduced access to recent state observations, which can be crucial for reacting to unexpected dynamics arising from modeling errors or environmental stochasticity. This tradeoff raises a crucial question: How can we preserve the strengths of action chunking in long-term consistency without suffering from its limitations in short-term reactivity?
To this end, we introduce Bidirectional Decoding (BID), an inference algorithm that bridges action chunking with closed-loop adaptation. Our main idea is to sample multiple predictions at each time step and search for the most desirable one. Specifically, BID operates on two decoding criteria: (i) backward coherence, which favors samples that are close to the sequence selected at the previous step; (ii) forward contrast, which favors samples that are close to the output of a stronger policy and distant from those of a weaker one. As illustrated in Fig. 1, BID updates the chunk of future actions based on the previous strategy, promoting temporal consistency over extended periods while remaining reactive to unexpected changes.
@@ 3.1 Preliminaries
Notably, Zhao et al. (2023b) hypothesize that action chunking mitigates compounding errors, but it is unclear why this would hold when deviations within a chunk cannot be corrected before replanning. Alternatively, Chi et al. (2023) draws parallels to model predictive control (MPC) (Borrelli et al., 2017; Löfberg, 2012), emphasizing its role in improving consistency and planning. However, unlike MPC, which typically uses short action horizons (e.g., $h=1$), recent imitation learning methods employ much longer action horizons, such as a substantial portion of the predicted sequence (e.g., $h=8$ in Chi et al. (2023)) or even the entire prediction length (e.g., $h\geq 50$ in Zhao et al. (2023b); Black et al. (2024)). The lack of a clear understanding of action chunking hinders its effective use across tasks and environments. We next explicitly analyze the strengths and weaknesses of action chunking, with particular attention to the choice of action horizon $h$.
@@ 3.2 Analysis
The divergence between a learned policy and an expert policy is attributed to two factors: (i) the importance of unobserved states in predicting the current action, and (ii) the difficulty of inferring unobserved states based on the available information.
To more clearly see the influence of action horizon on these factors, we next compare the performance of two policies that have the same context lengths but different action horizons, $\pi_{h}\coloneqq\pi_{(c,h)}(a_{t}|s_{t-h-c:t-h},a_{t-h:t-1})$ and $\pi_{h+d}\coloneqq\pi_{(c,h+d)}(a_{t}|s_{t-h-d-c:t-h-d},a_{t-h-d:t-1})$, where $d>0$ is the extended action horizon. As illustrated in Fig. 3, each policy has access to unique information that is unavailable to the other. $\pi_{h}$ observes some recent states, where $\pi_{h+d}$ is only aware of the executed actions. On the other hand, $\pi_{h+d}$ has access to some earlier states and actions, which precede all information available to $\pi_{h}$. We characterize the importance of observations as follows (formal definitions in Sec. E.2):
@@
Hence, we denote the observation advantage that $\pi_{h}$ gains from the observed recent states by $\alpha_{f}$ and the maximum inference disadvantage it incurs from the earlier unobserved states by $\epsilon_{b}$, whereas $\pi_{h+d}$ conversely gains $\alpha_{b}$ but incurs $\epsilon_{f}$.
Let $P(s_{t+1}\mid s_{t},a_{t})$ denote the true transition probabilities and $\hat{P}(s_{t+1}\mid s_{t},a_{t})$ the implicit dynamics model learned by the policy. The difficulty of inferring an unobserved state hinges on both the relevant observations and environmental stochasticity, which we quantify as follows.
Definition (Forward Inference). Let $P_{f}(t)\coloneqq P(S_{t}=g_{t}|S_{t-1}=g_{t-1},A_{t-1}=a_{t-1})$, where $g_{t}$ and $g_{t-1}$ are the ground truth states at time $t$ and $t-1$ respectively in a deterministic environment. In deterministic environments, $P_{f}(t)=1$; whereas in stochastic settings, $P_{f}(t)$ is smaller. The prediction error is characterized by $\delta_{f}(t)=\hat{P}(S_{t}=g_{t}|S_{t-1}=g_{t-1},a_{t-1})-P(S_{t}=g_{t}|S_{t-1}=g_{t-1},a_{t-1})$.
Definition (Backward Inference). Similarly, let $P_{b}(t)\coloneqq P(S_{t}=g_{t}|S_{t+1}=g_{t+1})$ where $g_{t}$ and $g_{t+1}$ are the ground truth states in the deterministic environment at time $t$ and $t+1$, respectively. Since $P_{b}(t)$ is not conditioned on any action, it has higher entropy in general. In stochastic environments, $P_{b}(t)$ is small. Furthermore, let $\delta_{b}(t)=\hat{P}(S_{t}=g_{t}|S_{t+1}=g_{t+1})-P(S_{t}=g_{t}|S_{t+1}=g_{t+1})$.
We refer to the variance of the distribution ${P}(a_{t^{\prime}-1},\ldots,a_{t^{\prime}-d}\mid s_{t^{\prime}})$ under expert policy as the diversity of past strategies up to time $t^{\prime}$. Given that the forward inference is generally easier than the backward inference, the performances of $\pi_{h}$ and $\pi_{h+d}$ differ as follows (proofs are deferred to Appendix E):
@@
Let $\mathcal{L}$ be a non-linear and non-negative convex function measuring the prediction error with respect to demonstrations. Let $C\coloneqq\{a_{t-h:t-1}\}\cup\mathcal{S}^{+}$ where $\mathcal{S}^{+}$ are the common states that both $\pi_{h}$ and $\pi_{h+d}$ observe. For the ease of notation, let $\tau_{f}=\{t-h-d+1:t-h\}$ and let $\tau_{b}=\{t-h-d-c:t-h-c-1\}$. Then, the difference in the expected loss between the $(c,h+d)$-policy and the $(c,h)$-policy, $\Delta_{d}\coloneqq\min_{\pi_{h+d}}\mathbb{E}\left[\mathcal{L}(\pi_{h+d},\pi^{*})|C\right]-\min_{\pi_{h}}\mathbb{E}\left[\mathcal{L}(\pi_{h},\pi^{*})|C\right]$, is bounded as:
$\alpha_{f}-\epsilon_{b}\left(1-\prod_{\tau\in\tau_{b}}P_{b}(\tau)(P_{b}(\tau)+\delta_{b}(\tau))\right)\leq\Delta_{d}\leq\epsilon_{f}\left(1-\prod_{\tau\in\tau_{f}}P_{f}(\tau)(P_{f}(\tau)+\delta_{f}(\tau))\right)-\alpha_{b}.$ (2)
@@
Remark 1. Eq. 2 provides a general comparison of the performance of the two policies. Intuitively, the advantage of each policy stems from the additional information it has access to (i.e. $\alpha_{f}$ for $\pi_{h}$ and $\alpha_{b}$ for $\pi_{h+d}$) while the disadvantage is bounded by the maximum divergence arising from inferring missing information incorrectly (i.e. $\epsilon_{b}$ and $\epsilon_{f}$ scaled by the maximum probability of incorrect inference).
We next examine the implications of Proposition 1.
Corollary 2 (Consistency).
Suppose the train and test environments are identical and deterministic. Suppose $a_{t}$ is influenced by at least one state at time steps $\tau_{b}$ and let $\delta_{f}(t^{\prime})\approx 0$ for all $t^{\prime}\in\tau_{f}$. If the diversity in past strategies up to time $t-h-c$ is not 0, and $\epsilon_{f}$ is finite, then
$\min_{\pi_{h+d}}\mathbb{E}\left[\mathcal{L}(\pi_{h+d},\pi^{*})|C\right]<$ $\min_{\pi_{h}}\mathbb{E}\left[\mathcal{L}(\pi_{h},\pi^{*})|C\right].$ (3)
@@
Remark 2. In deterministic environments, while both policies need to infer the same number of unobserved states, $\pi_{h+d}$ benefits from conditioning on additional actions, which may significantly aid in inferring the corresponding states through its action chunk. However, this is only true if the implicit dynamics model learned by the policy is approximately accurate and the maximum errors $\epsilon_{f}$ arising from inferring these states are bounded. Note that $\pi_{h+d}$ performs better as we increase the diversity of strategies present in the dataset i.e., as the distribution of actions taken by the expert prior to the context of the short horizon policy becomes higher variance.
Corollary 3 (Reactivity).
Suppose $P_{f}(t^{\prime})$ is small or both $P_{f}(t^{\prime})$ and $\left|\delta_{f}(t^{\prime})\right|$ are large for all $t^{\prime}\in\tau_{f}$. If temporal dependency decreases over time such that $\epsilon_{b}$ is small and $a_{t}$ is influenced by at least one state at time steps in $\tau_{f}$, then
$\min_{\pi_{h+d}}\mathbb{E}\left[\mathcal{L}(\pi_{h+d},\pi^{*})|C\right]>$ $\min_{\pi_{h}}\mathbb{E}\left[\mathcal{L}(\pi_{h},\pi^{*})|C\right].$ (4)
Remark 3. Recent states are often more important to decision making and, therefore, the disadvantages of action chunking become pronounced when inferring these recent states becomes difficult. This can happen if the test environment is stochastic. This can also happen if the test environment is deterministic as long as the implicit dynamics model is inaccurate - either due to a distribution shift or because of learning difficulty.
@@
In summary, depending on the experimental conditions, action chunking may have varying effects on the learned policy. On the one hand, it benefits the modeling of temporal dependencies in the demonstrations, due to extended access to actions executed at past time steps. On the other hand, it hinders reactions to unexpected dynamics, due to reduced access to state observations at recent time steps. As a result, there is no universally optimal choice of action horizon across all conditions. When both temporal dependencies and prediction errors are significant, tuning the action horizon entails an inherent trade-off between the two opposing factors.
@@ 4 Method: Bidirectional Decoding
As analyzed above, action chunking improves long-term consistency but sacrifices short-term reactivity. In this section, we propose to address this tradeoff by bridging action chunking with closed-loop adaptation. We will first outline the general framework in Sec. 4.1 and then describe two specific criteria in Sec. 4.2.
Figure 4: Illustration of bidirectional decoding.
Algorithm 1 Bidirectional Decoding
2:  Compute the backward loss $\mathcal{L}_{B}$ for each sample
4:  Compute the forward loss $\mathcal{L}_{F}$ for each sample
5:  Select $a^{*}\in\mathcal{A}$ that minimizes the total loss
6:  Update decision memory $\hat{a}\leftarrow a^{*}$
@@ 4.2 Bidirectional Criteria
Fig. 4 illustrates the combined effects of the backward coherence and forward contrast criteria on sample selection. Since samples in $\mathcal{A}^{+}$ and $\mathcal{A}^{-}$ are not all subject to the same strategy, we trim each set by removing samples that deviate significantly from the previous decision. This is achieved by summing over the $K$ smallest distance values for in the positive and negative sets in Eq. 7. The full process of our decoding method is outlined in Algorithm 1. Since all steps in BID can be computed in parallel, the overall computational overhead remains modest on modern GPU devices. Please refer to Appendix B for additional discussions about our decoding method.
5 Experiments
@@ 5.2.1 Comparison with Existing Inference Methods
We evaluate each method for $100$ episodes and average the results across three random seeds. Please refer to Appendix C for implementation details.
@@ 5.3 Real-world Experiments with Dynamic Objects
Setup. We consider two pick-and-place tasks, where the target object undergoes unexpected movement during evaluation. In the dynamic placing task, a Franka Panda robot is required to deliver an object into a moving cup held by a human subject. In the dynamic picking task, a UR5 robot is required to grasp a moving cup pulled by a string and place it onto a nearby static saucer. In both tasks, we evaluate the performance of BID applied to pre-trained diffusion policies. For further experimental details, please refer to Sec. C.2.